OpenStudy (anonymous):

3x/4+3=27 still dont get these

OpenStudy (anonymous):

32

OpenStudy (king):

32

OpenStudy (king):

3x+12/4=27 3x=96 x=32

OpenStudy (anonymous):

babs ill go through it step by step if u need

OpenStudy (anonymous):

\[(3x/4)+3=27\] \[(3x/4)=27-3\] \[(3x/4)=24\] \[3x=4*24\] \[x=32\]

OpenStudy (anonymous):

thanks having a hard time understanding lineat eguations it aint my best math prob

OpenStudy (anonymous):

3x/4+3=27 *4 *4 (Multiply each side by its denominator, in order to get rid of the fraction) 3x + 3 = 108 - 3 -3 (Subtract 3 from each side) 3x = 105 /3 /3 (Divide it by three) x = 35 Remember PEMDAS P =parenthesis E = exponent M = multiply D = divide A = add S = subtract

OpenStudy (anonymous):

\[(3x/4) + 3 = 27\] you want to get x by its self x = something in order to do that you subtract 3 from both sides of the equation \[(3x/4) + 3 -3 = 27 -3 \] \[3x/4 = 24\] now you multiply both sides by 4 in order to make it 3x=(27*4) then divide both sides by 3 and it be comes x=(27*4)/3 = 32

OpenStudy (anonymous):

you do the inverse of what is done to the x side of equation to cancel it out of that side and in order to make it so you dont change the equation what ever you do to one side you have to balance it by doing the same to the right side of the equation.

OpenStudy (anonymous):

lol thanks somebody finally exsplains it all to me cuz im like failing them with a 15% and it aint good

OpenStudy (anonymous):

you want to get x by it's self and you have to keep the equation balance think of it as a scale balance

OpenStudy (anonymous):

to get x by its self you need to subtract from it if something is added divide if its multiplied by something or the inverse

OpenStudy (anonymous):

post another one that u make up and ill explain that one to u try to stump me :-)

OpenStudy (anonymous):

no stumping u online math it suxs